Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
Answer by stanbon(75887)

Sally bought three chocolate bars and a pack of gum and paid $1.75. Jake bought two chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum.
----------------
1st: 3c+g=175
2nd: 2c+4g=200
--------
Multiply 1st by 2 and 2nd by 3 to get:
3rd: 6c+2g=350
4th: 6c+12g=600
-------------
Subtract 3rd from 4th to get:
10g=250
g=25
----------
Substitute into 2nd to solve for c:
2c+4*25=200
2c=100
c=50
-------------
# of chocolate bars = 50
# of packs of gum = 25
------
Cheers,
Stan H